OpenStudy (anonymous):

The volume of the mixture was 50 mL. Revising the above equation, she obtained? student studied the clock reaction described in this experiment. She set up Reaction Mixture 2 by mixing 20 mL 0.010 M KI, 10 mL .001 M Na2S2O3, 10 mL .040 M KBrO3, and 10 mL 0.10 M HCl using the procedure given. It took 45 seconds for the color to turn to blue. a. She found the concentration of each reactant in the reacting mixture by realizing that the number of moles of each reactant did not change when that reactant was mixed with the others, but that its concentration did.

OpenStudy (anonymous):

For any reactant A, no. moles A=MA Stock x VStock=M A Mixturex V Mixture The volume of the mixture was 50 mL. Revising the above equation, she obtained M A Mixture= M A stockX (VStock (mL))/ 50 mL Find the concentrations of each reactant by using the above equation. (I-)__________ M; (BrO3-) _____________M; (H+) ____________M b. what is the relative rate of the reaction (1000/t)? _________________ c. Knowing the relative rate of the reaction for mixture 2 and the concentrations of I-, BrO3-, and H+ in that mixture, she was able to set up equation 5 for the relative rate of the reaction, The only quantities that remained unknown were k,m,n, and p. Set up Equation 5 as she did, presuming she did it properly. 2. For reaction Mixture 1 the student found that 85 seconds were required. On dividing Equation 5 for Reaction Mixture 1 by Equation 5 for Reaction Mixture 2, and after canceling out the common terms (k, terms in (BrO3-) and (H+)), she got the following equation, 11.8/22=(.oo2o/.0040)m= (1/2)m recongnizing that 11.8/22 is about equal to 1/2, she obtained an approximate value for m. What is that value? m=_________________ By taking logarithms of both sides of the equation, she got an exact value for m. What was that value? m=_________________

OpenStudy (anonymous):

sorry no the student studied the clock reaction described in this experiment. She set up Reaction Mixture 4 by mixing 10 mL 0.010 M KI, 10 mL .001 M Na2S2O3, 10 mL .040 M KBrO3, and 10 mL 0.10 M HCl using the procedure given. It took 21 seconds for the color to turn to blue.

OpenStudy (anonymous):

also: 11.8/22=(.oo2o/.0040)m= (1/2)m recongnizing that 11.8/48 is about equal to 1/2, she obtained an approximate value for m. What is that value? m=_________________ By taking logarithms of both sides of the equation, she got an exact value for m. What was that value? m=_________________